OpenStudy (anonymous):

Help, Fan + medal!!!! Solve the equation -2/5n = -30

OpenStudy (anonymous):

A. n = –75 B. n = –12 C. n = 12 D. n = 75

OpenStudy (imstuck):

Solving for n...

OpenStudy (imstuck):

give me a sec to do it and then I'll help you, ok?

OpenStudy (anonymous):

K

OpenStudy (here_to_help15):

Divide both sides

OpenStudy (imstuck):

Is it this:\[-\frac{ 2 }{ 5n }=-30\]or\[-\frac{ 2 }{ 5 }n=-30\]

OpenStudy (here_to_help15):

-2n/5=30

OpenStudy (anonymous):

\[-\frac{ 2 }{ 5 }n = 30\]

OpenStudy (here_to_help15):

the answer would be -75

OpenStudy (imstuck):

ok, then the first thing you have to do is get rid of the 5 in the denominator by multiplying both sides by 5.

OpenStudy (imstuck):

-2n=150

OpenStudy (here_to_help15):

like i said -75

OpenStudy (imstuck):

now divide 150 by -2 and what do you get for n?

OpenStudy (here_to_help15):

oops srry for interupting

OpenStudy (anonymous):

-75

OpenStudy (anonymous):

Ok, and I have another question

OpenStudy (imstuck):

When you do order of operations in math, you first do parenthesis (but we don't have any here so..) then you deal with the multiplication and division, then the subtraction and addition. That also means that you undo those operations first. Which is why we got rid of the 5 in the denominator first.

OpenStudy (imstuck):

What is your other question?

OpenStudy (anonymous):

Joby's plant was 6 1/3 inches tall. Pat's was 4 1/2 inches tall. How much taller was Joby's plant than Pat's?

OpenStudy (imstuck):

Subtract pat's plant height from jody's plant height. Get them into improper fractions and then find the common denominator. Like this, one thing at a time...

OpenStudy (imstuck):